Calculate the concentration of hydrated copper sulfate, CuSO4.5H2O in a solution originally prepared by dissolving 0.10g of the solid in water and making up to 50cm3 with water?
1
Expert's answer
2020-07-07T14:21:40-0400

MOLAR MASS OF CUSO4.5H2O = 249.5 gm /mol ..


given mass of CUSO4.5H2O = 0.10 g .


So , moles of CUSO4.5H2O =( 0.10 / 249.5 ) mole .


volume of water = 50 cm3 .

mass of water = 50 gm ( density of water = 1 g / ml .) .



concentration of solution in terms of molality


=


moles of solute (CUSO4.5H2O ) / mass of solvent in kg ( water).




So , Molality = (0.10 /249 .5) mole . / (50 /1000 ) kg .


So , molality = ( 0.10 * 1000 ) / ( 249.5 * 50) mol /kg . = 0.008 m .